OpenStudy (anonymous):

Solve the given system, using the substitution method. y = 4x – 6 8x – 2y = 14 A. (14, 12) B. (12, 14) C. There are an infinite number of solutions. D. There is no solution.

OpenStudy (anonymous):

@samsterz

OpenStudy (anonymous):

hello??

OpenStudy (anonymous):

8x-2(4x-6)=14 8x-8x-6=14 0-6=14 -6 doesn't = 14 no solution

OpenStudy (anonymous):

Solve, using the substitution method. 4x + 2y = 11 x – 2 = – 2y

OpenStudy (anonymous):

A. 3, -1/2 B. 0, -2 C. 0, -11/2 D. 7, -5/2

OpenStudy (anonymous):

i think its a